3. You have discovered that a single gene controls the date when monarch butterflies break diapause (=become reproductively active) in Mexico. You were extremely busy this spring, and extensively sampled one overwintering site to find that 1751 monarchs broke diapause in early February, 424 broke diapause in late February, and 710 broke diapause in early March. You know there are two alleles for this gene (D, d), and that heterozygotes (Dd) break diapause in late February, while DD break diapause in early February. Showing your work, answer the following questions:

a. What is the observed frequency of the allele coding for early emergence (early February)?

b. Does this gene appear to be in Hardy-Weinberg equilibrium in this population?
Answer #1

Answer:

  • monarch 1600= early feb(MM)- alleles: 3200,
  • 400= late feb,(Mm)= total allele 800,so 400 early feb, and 400 early march as codominant
  • 700 early march(mm), 1400 alleles, total num of alleles 5400; 3600/5400=0.66(p)= early feb obs frequency;1800/5400=0.33(q)=early=obs freq of allele in early march.
  • to see H.W law, p^2+2pq+q^2=1; ..0.4356+.4356+.1089= 1, hence proved population is in H.W equilibrium

Please Rate My Answer.......Thank.......u...